# Plotting correlated data

### v1

@@
Abstract
A very common task in data visualization is to plot many data points with some measured y-value as a function of fixed x-values. Uncertainties on the y-values are typically presented as vertical error bars that represent either a Frequentist confidence interval or Bayesian credible interval for each data point. Most of the time, these error bars represent a 68% confidence/credibility level, which leads to the intuition that a model fits the data reasonably well if its prediction lies within the error bars of roughly two thirds of the data points. Unfortunately, this and other intuitions no longer work when the uncertainties of the data points are correlated. If the error bars only show the square root of diagonal elements of some covariance matrix with non-negligible off-diagonal elements, we simply do not have enough information in the plot to judge whether a drawn model line agrees well with the data or not. In this paper we will demonstrate this problem and discuss ways to add more information to the plots to make it easier to judge the agreement between the data and some model prediction in the plot, as well as glean some insight where the model might be deficient. This is done by explicitly showing the contribution of the first principal component of the uncertainties, and by displaying the conditional uncertainties of all data points.
I Introduction
A very common task in data visualization is to plot many data points with some measured y-value as a function of fixed x-values. Uncertainties on the y-values are typically presented as vertical error bars that represent either a Frequentist confidence interval or Bayesian credible interval for each data point. Most of the time, these error bars represent a 68% confidence/credibility level, which leads to the intuition that a model fits the data reasonably well if its prediction lies within the error bars of roughly two thirds of the data points. Unfortunately, this and other intuitions no longer work when the uncertainties of the data points are correlated. If the error bars only show the square root of diagonal elements of some covariance matrix with non-negligible off-diagonal elements, we simply do not have enough information in the plot to judge whether a drawn model line agrees well with the data or not.
@@ I Introduction
It is clear that just plotting the usual error bars is not sufficient to judge whether a model prediction matches the data or not. In such cases, at the very least, the squared M-distance of the model and the “degrees of freedom” need to be included in the plots to convey the actual goodness of fit. Alternatively, the corresponding p-value can be shown directly. Only showing the “reduced chi-squared”, where the squared M-distance is divided by the degrees of freedom is not sufficient, since it does not uniquely correspond to a p-value.
II Plotting the correlation matrix
@@
Figure 2: Examples of colour maps used to plot correlation matrices and their names in the Matplotlib library[1]. Divergent maps like LABEL:sub@fig:coolwarm show a clear distinction between positive and negative values, but they do not work without the colour information LABEL:sub@fig:coolwarm-gray. Perceptually uniform sequential colour maps like LABEL:sub@fig:cividis and LABEL:sub@fig:gray work well even in monochrome, but they make it harder to distinguish small positive from small negative values. The correlation matrix is taken from the $\delta p_{T}$ result in [2, 3].
Hinton diagrams – originally introduced to show weights in a neural network[4] – can overcome this issue. In these, the absolute value of the matrix elements is not represented by the hue or lightness of the colour, but by the area of a symbol drawn at each position corresponding to a matrix element. The background colour is taken from the middle of the colour scale, and the colour of the symbols from both ends. One indicates positive values and one negative (see Figure 3). Small bright dots are very easily distinguished from small dark dots. When using circles as symbols, Hinton diagrams work well with surprisingly big matrices. As the circles get smaller and smaller, the overall impression of the diagram becomes like a regular 2D histogram where the colour determines the element value, since it effectively turns into a halftone image[5].
@@
Figure 3: Examples of Hinton plots and the names of the used colour maps in the Matplotlib library[1]. In these, the absolute value of the matrix elements is reflected as the area of the circles, while the colour of the circles represents the sign of the value. Positive values are easily distinguishable from negative values even for small absolute values and a complete lack of colour information. The correlation matrix is taken from the $\delta p_{T}$ result in [2, 3].
@@
The presence of the correlation matrix makes it possible to understand why a model fits better or worse to the data than expected, but spreading the information over two separate plots is not ideal. Especially with higher number of data points it can become quite bothersome to try to assign the right matrix element to any given combination of data point, and keep all that information in mind while scanning the actual data plot for differences between the data and the models. It would be more convenient to embed some information about the correlations in the main plot itself.
@@
III Showing the correlation between neighbouring data points
@@
(a) The example data with correlation lines indicating the correlation between neighbouring data points. The position where the lines connect to the error bars relative to the length of the error bars corresponds to the absolute value of the correlation coefficient between the two bins. Positive correlation is indicated by lines both connecting on the same side of the data points, anticorrelation by crossing lines which connect to opposing sides. If there is no correlation, there will be just a single line connecting the data points directly.
(b) The example data shown as a principal component plot. The hatched areas indicate that the principal component of the covariance strongly correlates the last two bins and anticorrelates them with the first bin. When this component is reduced, the remaining covariance has positive correlations between all neighbours as indicated by the correlation lines. The prediction of model M2 does not align with any of those correlations. See Figure 6 for a detailed explanation of the different features in the plot.
Figure 5:
Aside from giving an impression of the sign and magnitude of the correlation between the data points, some quantitative things can be read off these lines. Assuming multivariate normal uncertainties, the expected values of the bins change when fixing one of the neighbouring bins according to[6, P88]
@@
One can also rearrange this equation to yield the law of total variance for a data point[6, P292]:
@@
IV Showing the dominant principal component of the uncertainty
@@
A general way of identifying the most important components of a multivariate distribution is the Principal Component Analysis (PCA, see e.g. [7]). The principal components are the $N$ unit-length eigenvectors of the correlation matrix $\bm{u}_{i}$, ordered by their corresponding eigenvalues $\lambda_{i}$ from largest to smallest. If the uncertainties are uncorrelated, all eigenvalues will be 1. In the presence of correlations, the eigenvalues will no longer be equal, but by construction they will always sum up to $N$. The eigenvalues are a measure of how much each component contributes to the total variance in the uncertainty.
@@
When plotting the data with error bars corresponding to both $C$ and $K$, the difference between the error bars show the contribution that has been removed. Importantly, the removed component stems from a single, 100% correlated variation in the data along the direction of $W\bm{u}_{1}$. So these differences have a defined direction in the N-dimensional data space. This direction can be indicated by hatching the area between the full and remaining covariance errors differently depending on whether the error bars are on the same side as the corresponding component of $\bm{u}_{i}$ or the opposite. This is done in Figure 5b, and the concept is explained and sketched out in more detail in Figure 6.
Figure 6: The example data shown in all pairwise 2D projections. The solid ellipses show the original covariance, i.e. where the M-distance of the marginalised uncertainties is 1. These ellipses determine the total length of the error bars, which correspond to the square root of the diagonal elements of the covariance matrix. The contribution of the first principal component of the covariance is reduced, and the remaining covariance is plotted as dotted lines. The diagonal elements of this remaining covariance determine the inner edges of the hatched areas. So the hatched areas show uncertainty that is caused by the first principal component. The type of hatching indicates whether the area is in the positive or negative direction of that component. The dashed ellipses show the conditional covariance of the full uncertainties when all other $\Delta y$ are assumed to be 0. I.e. it is a slice of the N-dimensional ellipsoid rather than a projection. Where these ellipses intersect the axes indicates the conditional uncertainty of that data point, assuming all others are fixed. This determines the position of the middle points of the triangles. If a model prediction resides on the side of the data with the same hatching type, it benefits from the largest uncertainties, indicated by the outer edges of the hatched areas. If it is not clearly in either of those directions, it can be compared to the remaining covariance, indicated by the inner edges of the hatched areas. The remaining covariance matrix still has correlations in it and for neighbouring bins those are indicated by the correlation lines in Figure 5b (not shown here).
@@
A conservative choice is to scale the contribution of the first principal component such that it is equal to the contribution of the second component. To do this, one chooses $s^{2}=(1-\lambda_{2}/\lambda_{1})$. The choice of the “target” eigenvalue is somewhat arbitrary, though, and the ideal value depends on the plotted data. A reasonable starting point is to aim for the median value: $s^{2}=(1-\lambda_{m}/\lambda_{1})$. If one wants to get the maximum effect in the plot, and assuming that the original covariance matrix is not rank deficient, one can also aim for the smallest eigenvalue: $s^{2}=(1-\lambda_{N}/\lambda_{1})$. Figures 5b and 6 use the median target value, while figures 7 and 8 use the smallest.
@@
Of course it is also possible to show not just the first principal component, but the first $N_{C}$ components. In this case, there will be $N_{C}$ scaling factors, which should be chosen so the target eigenvalue is the same for all of them. $K$ is now the remaining covariance after subtracting the contributions of the $N_{C}$ components, and hatched areas are shown successively by in turn adding back the contributions from smallest to largest eigenvalue, and filling the difference between the new and previous error bars. Such plots can get very busy, though, and beyond the first or maybe second component, the readability drops sharply. Examples of this are shown in appendix A.
V Conditional uncertainties
@@
These conditional uncertainties of the full covariance $V$ are shown as the inner points of the triangles in figures 5b and 6.
The conditional uncertainties represent slices through the parameter space (fixing other variables), as opposed to marginal uncertainties, which represent projections (ignoring other variables). If the covariance matrix in question is uncorrelated, the conditional and marginal uncertainties are identical. If the covariance is considerably more constrained in one directions compared to the others, i.e. if one eigenvalue of the correlation matrix is much smaller than the others, then the conditional uncertainties will correspond to the uncertainty in the direction of this strongest constraint. The inclusion of the conditional uncertainties can thus give another indication of correlations and is complementary to the show of principal components.
VI Combined effect
Compared to the current default style in Figure 1, plots in the style of Figure 5b should enable authors to convey much more useful information about the shape of uncertainties in their data. Naturally, this comes at the “cost” of making the plots more information dense and thus a bit harder to parse. But the plots only contain additional information, and when in doubt, it is always possible to completely ignore the additional bits and only read off the outer limits of the hatched rectangles. These correspond exactly to the marginal error bars shown in the default plot style. For further interpretation of visual comparisons to model predictions, one can employ these rules of thumb:
@@ VI Combined effect
VII A real-world example
@@
Figure 7 shows the result of the $\delta p_{T}$ cross-section measurement from [2, 3] in the new principal component plot style. The correlation matrix of this result is shown in figures 2 and 3. The hatched areas clearly show the considerable anticorrelations between the second, third, and fourth data points, suggesting that the “dip” in that region is not physical, but a statistical fluctuation.
@@
Figure 7: The $\delta p_{T}$ result from [2, 3] shown as a principal component plot. The nominal Monte Carlo model was provided by the original authors of the measurement. The strong anticorrelation of the second, third, and fourth bins are clearly visible in the first principal component. The regularised result is on the “right” side of all three data points, while the nominal model is not. The model thus “sees” only the remaining uncertainties without the first component. The correlation lines for the remaining covariance were omitted in this plot, since the remaining correlation was negligible and only reduced the readability of the data.
@@
The error bars for the last data point in Figure 7 are too small to be seen, which makes it impossible to judge how much that bin contributes to the bad agreement between the data and the nominal model. This can be easily fixed by plotting the data to model ratio, as done in Figure 8. It also includes the local gradient of the Mahalanobis distance between the model and the data, as proposed in [8]. Together, this makes it clear that while the dip in the third bin might be visually the most significant deviation of the model from the data in Figure 7, the discrepancy is actually more driven by the first, and the two last bins.
@@
Figure 8: The same data as in Figure 7 shown as a principal component plot combined with the model ratio and local gradient of the M-distance as proposed in [8]. The gradient was scaled to minimise the squared M-distance at its endpoint, which is shown in the legend. Compared to Figure 7, the anticorrelations of the first principal component are now better visible in the first 5 data points. Due to these correlations, the gradient is not pointing directly at the data points. For one data point it is even pointing away from the data, meaning the model fit to the data is improved by moving it away from that data point.
VIII Conclusions
@@
The practice of plotting data points with only their marginal uncertainties can produce very misleading plots if there are considerable correlations in the covariance of the data. The proposed plotting styles can add additional information to plots, that make it easier to spot effects of such correlations. The correlation lines are an easy way to include correlations between neighbouring data points. In case there are no correlations, they will behave like a simple line connecting the data points. The principal component plots can add information off the strength and direction of the largest principal components of the covariance. The inclusion of conditional uncertainties in the plot can indicate the presence of strong constraints in the uncertainty that are not reflected in the size of the marginal error bars. In case there are no correlations, they will revert to look like the usual “marginal error only” plots. In either case, the correlation information is purely additive, and can easily be ignored, yielding the same information that the classical plots provide.
@@
The plotting styles proposed in this paper are implemented in the “NuStatTools” Python package, and available online[9]. The plots in this document were generated with version 0.8.0 of the package. The Jupyter Notebook with the actual plotting commands for the artificial examples is available as supplementary material. The data shown in section VII is available in their data release[3].

### v2

@@
Plotting Correlated Data
Lukas Koch
Johannes Gutenberg University Mainz
lukas.koch@uni-mainz.de
[…]
Institute of Physics – ETAP
Johannes Gutenberg University Mainz
Staudingerweg 7, 55128 Mainz
E-mail:
1 Introduction
The visualisation of data with uncertainties is non-trivial and an active field of research (see e.g. Kamal et al., (2021) for an overview). The choice of a “best” visualisation technique can depend on the type of data, as well as the nature of the uncertainties, and the intended audience of the visualisation. In this work we will focus on a common task in most quantitative sciences: the plotting of many data points with some measured $y$-values as a function of fixed $x$-values, intended for consumption by other researchers.
In these cases, uncertainties on the $y$-values are typically presented as vertical error bars that represent either a Frequentist confidence interval or Bayesian credible interval for each data point. Most of the time, these error bars represent a 68% confidence/credibility level, which leads to the intuition that a model fits the data reasonably well if its prediction lies within the error bars of roughly two thirds of the data points. Unfortunately, this and other intuitions no longer work when the uncertainties of the data points are correlated. In those cases, the error bars only show the square root of the diagonal elements of the covariance matrix. If the covariance has non-negligible off-diagonal elements, we simply do not have enough information in the plot to judge whether a drawn model line agrees well with the data or not.
It is clear that just plotting the usual error bars is not sufficient to judge whether a model prediction matches the data or not. In such cases, at the very least, the squared M-distance of the model and the “degrees of freedom” need to be included in the plots to convey the actual goodness of fit. Alternatively, the corresponding $p$-value can be shown directly. Only showing the “reduced chi-squared”, where the squared M-distance is divided by the degrees of freedom, is not sufficient since it does not uniquely correspond to a $p$-value.
@@ 1 Introduction
2 Plotting the Correlation Matrix
@@
[…]
Hinton diagrams – originally introduced to show weights in a neural network (Hinton and Shallice,, 1991) – can overcome this issue. In these, the absolute value of the matrix elements is not represented by the hue or lightness of the colour, but by the area of a symbol drawn at each position corresponding to a matrix element. The background colour is taken from the middle of the colour scale, and the colour of the symbols from both ends. One indicates positive values and one negative (see Figure 3). Small bright dots are very easily distinguished from small dark dots. When using circles as symbols, Hinton diagrams work well with surprisingly big matrices. As the circles get smaller and smaller, the overall impression of the diagram becomes like a regular 2D histogram where the colour determines the element value, since it effectively turns into a halftone image (Campbell,, 2000).
@@
[…]
@@
The presence of the correlation matrix makes it possible to understand why a model fits better or worse to the data than expected, but spreading the information over two separate plots is not ideal. Especially with higher number of data points it can become quite bothersome to try to assign the right matrix element to any given combination of data points, while scanning the actual data plot for differences between the data and the models. It would be more convenient to embed some information about the correlations in the main plot itself.
@@
3 Correlations between Neighbouring Data Points
@@
These lines are similar in concept to parallel coordinates plots (see e.g. Moustafa, (2011)). But where the latter can be used to show arbitrary relations between dimensions of samples from a random distribution, the correlation lines show a single summary statistic of the relation between neighbouring dimensions in the whole distribution. For multivariate normal distributed data, the correlation lines convey all information with much less visual clutter than classical parallel coordinate plots, as shown in Figure 5b.
(a) Correlation lines
(b) Random throws
Figure 5: LABEL:sub@fig:corlines The example data with correlation lines indicating the correlation between neighbouring data points. The position where the correlation lines connect to the error bars relative to the length of the error bars corresponds to the absolute value of the correlation coefficient between the two bins. Positive correlation is indicated by lines both connecting on the same side of the data points, anticorrelation by crossing lines which connect to opposing sides. If there is no correlation, there will be just a single line connecting the data points directly. Compared to the plot of random samples from the covariance as shown in LABEL:sub@fig:parallel, the correlation lines introduce much less clutter.
Figure 6: The example data as a principal component plot. The hatched areas indicate that the principal component of the covariance strongly correlates the last two bins (as the hatch stiles match on the same sides of the data points) and anticorrelates them with the first bin (as the hatch styles match on opposing sides of the data points). When this component is reduced, the remaining covariance has positive correlations between all neighbours as indicated by the correlation lines. The prediction of model M2 does not align with any of those correlations. See Figure 7 for a detailed explanation of the different features in the plot.
Aside from giving an impression of the sign and magnitude of the correlation between the data points, some quantitative things can be read off the correlation lines. Assuming multivariate normal uncertainties, the expected values of the bins change when fixing one of the neighbouring bins according to (Soch et al.,, 2025, P88)
@@
One can also rearrange this equation to yield the law of total variance for a data point (Soch et al.,, 2025, P292):
@@
4 Showing the Dominant Principal Component
@@
A general way of identifying the most important components of a multivariate distribution is the Principal Component Analysis (PCA, see e.g. Bohm and Zech, (2017)). The principal components are the $N$ unit-length eigenvectors of the correlation matrix $\bm{u}_{i}$, ordered by their corresponding eigenvalues $\lambda_{i}$ from largest to smallest. If the uncertainties are uncorrelated, all eigenvalues will be 1. In the presence of correlations, the eigenvalues will no longer be equal, but by construction they will always sum up to $N$. The eigenvalues are a measure of how much each component contributes to the total variance in the uncertainty.
@@
When plotting the data with error bars corresponding to both $C$ and $K$, the difference between the error bars show the contribution that has been removed. Importantly, the removed component stems from a single, 100% correlated variation in the data along the direction of $W\bm{u}_{1}$. So these differences have a defined direction in the N-dimensional data space. This direction can be indicated by hatching the area between the full and remaining covariance errors differently depending on whether the error bars are on the same side as the corresponding component of $\bm{u}_{i}$ or the opposite. This is done in Figure 6, and the concept is explained and sketched out in more detail in Figure 7.
Figure 7: The example data shown in all pairwise 2D projections. The solid ellipses show the original covariance, i.e. where the M-distance of the marginalised uncertainties is 1. These ellipses determine the total length of the error bars, which correspond to the square root of the diagonal elements of the covariance matrix. The contribution of the first principal component of the covariance is reduced, and the remaining covariance is plotted as dotted lines. The diagonal elements of this remaining covariance determine the inner edges of the hatched areas. So the hatched areas show uncertainty that is caused by the first principal component. The type of hatching indicates whether the area is in the positive or negative direction of that component. The remaining covariance matrix still has correlations in it and for neighbouring bins those are indicated by the correlation lines in Figure 6 (not shown here). The dashed ellipses show the conditional covariance of the full uncertainties when all other $\Delta y$ are assumed to be 0. I.e. it is a slice of the N-dimensional ellipsoid rather than a projection. Where these ellipses intersect the axes indicates the conditional uncertainty of that data point, assuming all others are fixed. This determines the position of the middle points of the triangles.
@@
A conservative choice is to scale the contribution of the first principal component such that it is equal to the contribution of the second component. To do this, one chooses $s^{2}=(1-\lambda_{2}/\lambda_{1})$. The choice of the “target” eigenvalue is somewhat arbitrary, though, and the ideal value depends on the plotted data. A reasonable starting point is to aim for the median value: $s^{2}=(1-\lambda_{m}/\lambda_{1})$. If one wants to get the maximum effect in the plot, and assuming that the original covariance matrix is not rank deficient, one can also aim for the smallest eigenvalue: $s^{2}=(1-\lambda_{N}/\lambda_{1})$. Figures 6 and 7 use the median target value, while Figures 8 and 9 use the smallest.
@@
Of course, it is also possible to show not just the first principal component, but the first $N_{C}$ components. In this case, there will be $N_{C}$ scaling factors, which should be chosen so the target eigenvalue is the same for all of them. $K$ is now the remaining covariance after subtracting the contributions of the $N_{C}$ components. Hatched areas are then shown successively, by adding back the contributions from smallest to largest eigenvalue in turn, and filling the difference between the new and previous error bars. Such plots can get very busy, though, and beyond the first or maybe second component, the readability drops sharply. Examples of this are shown in appendix A.
5 Conditional Uncertainties
@@
These conditional uncertainties of the full covariance $V$ are shown as the inner points of the triangles in Figures 6 and 7.
The conditional uncertainties represent slices through the parameter space (fixing other variables), as opposed to marginal uncertainties, which represent projections (ignoring other variables). If the covariance matrix in question is uncorrelated, the conditional and marginal uncertainties are identical. If the covariance is considerably more constrained in one directions compared to the others, i.e. if one eigenvalue of the correlation matrix is much smaller than the others, then the conditional uncertainties will correspond to the uncertainty in the direction of this strongest constraint. The inclusion of the conditional uncertainties can thus give another indication of correlations and is complementary to the display of principal components.
6 Combined Effect
Compared to the current default style in Figure 1, plots in the style of Figure 6 should enable authors to convey much more useful information about the shape of uncertainties in their data. Naturally, this comes at the “cost” of making the plots more information dense and thus a bit harder to parse. But the plots only contain additional information, and when in doubt, it is always possible to completely ignore the additional bits and only read off the outer limits of the hatched rectangles. These correspond exactly to the marginal error bars shown in the default plot style. For further interpretation of visual comparisons to model predictions, one can employ these rules of thumb:
@@ 6 Combined Effect
7 A Real-World Example
@@
Figure 8 shows the result of the $\delta p_{T}$ cross-section measurement from Abe et al., 2018a ; Abe et al., 2018b in the new principal component plot style. The correlation matrix of this result is shown in Figures 2 and 3. The hatched areas clearly show the considerable anticorrelations between the second, third, and fourth data points, suggesting that the “dip” in that region is not physical, but a statistical fluctuation.
@@
Figure 8: The $\delta p_{T}$ result from Abe et al., 2018a ; Abe et al., 2018b shown as a principal component plot. The nominal Monte Carlo model was provided by the original authors of the measurement. The strong anticorrelation of the second, third, and fourth bins are clearly visible in the first principal component. The regularised result is on the “right” side of all three data points, while the nominal model is not. The model thus “sees” only the remaining uncertainties without the first component. The correlation lines for the remaining covariance were omitted in this plot, since the remaining correlation was negligible and only reduced the readability of the data.
@@
The error bars for the last data point in Figure 8 are too small to be seen, which makes it impossible to judge how much that bin contributes to the bad agreement between the data and the nominal model. This can be easily fixed by plotting the data to model ratio, as done in Figure 9. It also includes the local gradient of the Mahalanobis distance between the model and the data, as proposed in Koch, (2022). Together, this makes it clear that the discrepancy is actually more driven by the first and the two last bins, even if the dip in the third bin might be visually the most significant deviation of the model from the data in Figure 8,
@@
Figure 9: The same data as in Figure 8 shown as a principal component plot combined with the model ratio and local gradient of the M-distance as proposed in Koch, (2022). The gradient was scaled to minimise the squared M-distance at its endpoint, which is shown in the legend. Compared to Figure 8, the anticorrelations of the first principal component are now better visible in the first 5 data points. Due to these correlations, the gradient is not pointing directly at the data points. For one data point it is even pointing away from the data, meaning the model fit to the data is improved by moving it away from that data point.
This is just a single example and cannot guarantee that this type of plot will always be useful in conveying information about the correlation between data points. Every data set and every correlation structure is different, and an exhaustive analysis of when principal component plots make sense is beyond the scope of this paper.
8 Conclusions
The practice of plotting data points with only their marginal uncertainties can produce very misleading plots if there are considerable correlations in the covariance of the data. The proposed plotting styles can add additional information to plots that make it easier to spot effects of such correlations.
@@ 8 Conclusions
The correlation lines are an easy way to include correlations between neighbouring data points. In case there are no correlations, they will behave like a simple line connecting the data points. The principal component plots can add information off the strength and direction of the largest principal components of the covariance. The inclusion of conditional uncertainties in the plot can indicate the presence of strong constraints in the uncertainty that are not reflected in the size of the marginal error bars. In case there are no correlations, they will revert to look like the usual “marginal error only” plots. In either case, the correlation information is purely additive, and can easily be ignored, yielding the same information that the classical plots provide.
@@
As a rule of thumb, it is usually a good idea to provide a plot of the full correlation matrix in the form of a Hinton diagram. It contains the full correlation information between all dimensions of the data, even if it can be a bit hard to parse. Then, depending on the actual structure of the data and correlations, the correlation lines or principal component plots can be used to insert additional information into the data plot itself. The choice of which to use depends on the data, and it is difficult to make general judgements. One should try and check which plotting style adds the most information without overloading the plot. Again, as a rule of thumb, if there is a single dominating principal component, the PC plot will probably give good results. If the data points are not too “squished together” in the $x$-direction and correlations are mostly short range, the correlation lines will likely add useful information.
@@
[…]